1)find the least value of 6x2-22x+21/5x2-18x+17 for real x
2)solve for x (x2-4x)2 – (x-2)2 – 16 = 0
3)the cartesian product of A*A has 9 elements among which are found (-1,0) and (0,1). find the set A
4)Find the domain fo the function
f(x) = 1/sq root x-IxI.
these are cbse questions so pls give answers with steps
 

Grade:12

2 Answers

Vikas TU
7 years ago
 6x2-22x+21/5x2-18x+1
to minimise it maximise the denominator and minimise the numerator simultaneously.
Therefore,
N(x) =  6x2-22x+21
N(x)’ = 12x – 22 
N(x)’’ =  12 > 0 .i.e 12 x – 22 =0
x = 11/6 is the minima point.
D(x) = 5x2-18x+1
D(x)’ = 10x – 18
D(x)’’ = 10 > 0 
thus,
10 x – 18 = 0
x = 9/5 is the minima point but we want to maximis ethe denominator therefore, neglecting x = 9/5.
Put x = 11/6 in  6x2-22x+21/5x2-18x+1 that would be the minum vale of the expression.
Vikas TU
7 years ago
Dear raghav,
plz put your questions in sequence wise at once.
At one time post your question as a one time.
In that way it would be easy for us to answer sequentially.
Thankyou!